# 4.2cm4.2cm3.6cm3cm

### v1

@@ 2. Derivations on $\mathcal{L}(\Gamma)$
For all $\alpha\in\Gamma,i\in\mathbb{Z}$, let $D(H_{\alpha,i})=\phi_{\alpha,i}L_{\alpha,i}+\psi_{\alpha,i}H_{\alpha,i},\,\phi_{\alpha,i},\psi_{\alpha,i}\in\mathbb{F}[t^{1},t^{-1}].$
$D([L_{-\alpha,-i},H_{\alpha,i}])$ $=$ $D(-\alpha H_{0,0})$
$=$ $-\alpha(\phi_{0,0}L_{0,0}+\psi_{0,0}H_{0,0})$.
@@
$D([L_{-\alpha,-i},H_{\alpha,i}])$ $=$ $[D(L_{-\alpha,-i}),H_{\alpha,i}]+[L_{-\alpha,-i},D(H_{\alpha,i})]$
$=$ $[f_{-\alpha}L_{-\alpha,-i}+g_{-\alpha}H_{-\alpha,-i},H_{\alpha,i}]+[L_{-\alpha,-i},\phi_{\alpha,i}L_{\alpha,i}+\psi_{\alpha,i}H_{\alpha,i}]$
$=$ $-\alpha f_{-\alpha}H_{0,0}-2\alpha\phi_{\alpha,i}L_{0,0}-\alpha\phi_{\alpha,i}H_{0,0}$.
We have $\alpha(2\psi_{\alpha,i}-\psi_{0,0})L_{0,0}+\alpha(f_{-\alpha}+\phi_{\alpha,i}-\phi_{0,0})H_{0,0}=0$, thus
$\phi_{\alpha,i}=\frac{1}{2}\phi_{0,0},\,f_{-\alpha}+\psi_{\alpha,i}=\psi_{0,0},\,\forall 0\neq\alpha\in\Gamma.$
Since $H_{0,0}\in C(\mathcal{L}(\Gamma))$, then $D(H_{0,0})\in C(\mathcal{L}(\Gamma))$, thus $\phi_{0,0}=0$, so $\phi_{\alpha,i}=0,\,\forall\alpha\in\Gamma,i\in\mathbb{Z}$, we have
$D(H_{\alpha,i})$ $=$ $\psi_{\alpha,i}H_{\alpha,i}$
$=$ $(\psi_{0,0}-f_{-\alpha})H_{\alpha,i}$
$=$ $(f_{\alpha}+\psi_{0,0})H_{\alpha,i}$.
@@
$\mathrm{Der}\mathcal{L}(\Gamma)=\mathrm{ad}(\mathcal{L}(\Gamma))\bigoplus(\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}\oplus\mathcal{D}_{\mathbf{g}(\Gamma)}\oplus\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]}\oplus\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}})$.
@@
(i)$(\mathrm{Der}\mathcal{L}(\Gamma))_{0}=\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}+\mathcal{D}_{\mathbf{g}(\Gamma)}+\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]}+\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}}$.
In fact, by Lemma 3 and Lemma 4, for all $D\in\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}$, there exists $\rho\in\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt},f\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]),g\in\mathbf{g}(\Gamma),b\in\mathbb{F}[t^{1},t^{-1}]$, such that
$(D-D^{\rho})(L_{\alpha,i})$ $=$ $f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i}$,
$(D-D^{\rho})(H_{\alpha,i})$ $=$ $(f_{\alpha}+b)H_{\alpha,i}$.
On the other hand, it follows from the definition of $D_{f},D_{g}$ and $D_{b}$ that
$(D_{f}+D_{g}+D_{b})(L_{\alpha,i})$ $=$ $f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i}$,
$(D_{f}+D_{g}+D_{b})(H_{\alpha,i})$ $=$ $(f_{\alpha}+b)H_{\alpha,i}$.
Thus $D=D^{\rho}+D_{f}+D_{g}+D_{b}$.
(ii)$\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}+\mathcal{D}_{\mathbf{g}(\Gamma)}+\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]}+\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}}$ is a direct sum.
@@
$D=D_{\phi}+\sum_{j\in J}c_{j}D_{g_{j}}+D_{b}+D^{\rho}.$
$J$ is a finite index set, $\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]),\,g_{j}\in\mathbf{g}(\Gamma),\,b\in\mathbb{F}[t^{1},t^{-1}],\,\rho\in\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt},\,c_{j}\in\mathbb{F}$. If $D=0$, to prove that $D_{\phi}=\sum_{j\in J}c_{j}D_{g_{j}}=D_{b}=D^{\rho}=0$, we have
$D(L_{\alpha,i})$ $=$ $0=f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i},$
$D(H_{\alpha,i})$ $=$ $0=(f_{\alpha}+b)H_{\alpha,i}.$
@@
$(D_{\phi}+\sum_{j\in J}c_{j}D_{g_{j}}+D_{b}+D^{\rho})(L_{\alpha,i})$ $=$ 00,
$(D_{\phi}+\sum_{j\in J}c_{j}D_{g_{j}}+D_{b}+D^{\rho})(H_{\alpha,i})$ $=$ 00.
Then $D_{\phi}=\sum_{j\in J}c_{j}D_{g_{j}}=D_{b}=D^{\rho}=0$. ∎
@@ 3. $2$-local derivations on $\mathcal{L}(\Gamma)$
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$. For all $x,y\in\mathcal{L}(\Gamma)$, there exist derivations $\Delta_{x,y}$ on $\mathcal{L}(\Gamma)$, such that $\Delta(x)=\Delta_{x,y}(x)\,,\,\Delta(y)=\Delta_{x,y}(y)$, and it can be written as
$\Delta_{x,y}=\mathrm{ad}(\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ i\in\mathbb{Z}\end{subarray}}(a_{\alpha,i}(x,y)L_{\alpha,i}+b_{\alpha,i}(x,y)H_{\alpha,i}))+\lambda_{1}(x,y)D_{f}+\lambda_{2}(x,y)D_{g}+\lambda_{3}(x,y)D_{b}+\lambda_{4}(x,y)D^{\rho}.$
where $a_{\alpha,i},b_{\alpha,i},\lambda_{i}(i=1,2,3,4)$ are complex-valued functions on $\mathcal{L}(\Gamma)\times\mathcal{L}(\Gamma)$, $D_{f},D_{g},D_{b},D^{\rho}$ is given previous.
@@
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$. For any fixed $x\in\mathcal{L}(\Gamma)$, if $\Delta(L_{\alpha,i})=0$, $\forall\alpha\in\Gamma,i\in\mathbb{Z}$, then
$\Delta_{L_{\alpha,i},x}$ $=$ $\mathrm{ad}(\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}(a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha,j}+b_{0,j}(L_{\alpha,i},x)H_{0,j}))+(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)D_{f}$
$+(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)D_{g}+\lambda_{3}(L_{\alpha,i},x)D_{b}+(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)D^{\rho}$.
@@
If $\Delta(L_{\alpha,i})=0$, we have
$\Delta(L_{\alpha,i})$ $=$ $\Delta_{L_{\alpha,i},x}(L_{\alpha,i})$
$=$ $[\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}(a_{\beta,j}(L_{\alpha,i},x)L_{\beta,j}+b_{\beta,j}(L_{\alpha,i},x)H_{\beta,j}),L_{\alpha,i}]+\lambda_{1}(L_{\alpha,i},x)D_{f}(L_{\alpha,i})$
$+\lambda_{2}(L_{\alpha,i},x)D_{g}(L_{\alpha,i})+\lambda_{3}(L_{\alpha,i},x)D_{b}(L_{\alpha,i})+\lambda_{4}(L_{\alpha,i},x)D^{\rho}(L_{\alpha,i})$
$=$ $\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}((\beta-\alpha)a_{\beta,j}(L_{\alpha,i},x)L_{\alpha+\beta,i+j}+\beta b_{\beta,j}(L_{\alpha,i},x)H_{\alpha+\beta,i+j})$
$+f_{\alpha}\lambda_{1}(L_{\alpha,i},x)L_{\alpha,i}+g_{\alpha}\lambda_{2}(L_{\alpha,i},x)L_{\alpha,i}+i\lambda_{4}(L_{\alpha,i},x)L_{\alpha,i-1}$
$=$ 00.
Thus $a_{\beta,j}(L_{\alpha,i},x)=0\,(\alpha\neq\beta),\,b_{\beta,j}(L_{\alpha,i},x)=0\,(\beta\neq 0),\,\lambda_{1}(L_{0,i},x)=\lambda_{2}(L_{0,i},x)=\lambda_{4}(L_{\alpha,0},x)=0.$
@@
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$ such that $\Delta(L_{0,i})=\Delta(L_{1,j})=0$. Then $\Delta(L_{\alpha,k})=0,\,\forall\alpha\in\Gamma,\,k\in\mathbb{Z}$.
@@
Since $\Delta(L_{0,i})=\Delta(L_{1,j})=0$, we can assume
$\Delta_{L_{\gamma,k},x}$ $=$ $\mathrm{ad}(a_{\gamma,j}(L_{\gamma,k},x)L_{\gamma,j}+b_{0,j}(L_{\gamma,k},x)H_{0,j})+(1-\delta_{\gamma,0})\lambda_{1}(L_{\gamma,k},x)D_{f}$
$+(1-\delta_{\gamma,0})\lambda_{2}(L_{\gamma,k},x)D_{g}+\lambda_{3}(L_{\gamma,k},x)D_{b}+(1-\delta_{k,0})\lambda_{4}(L_{\gamma,k},x)D^{\rho},\,\gamma=0,1,\,\forall x\in\mathcal{L}(\Gamma)$.
Let $x=L_{\alpha,k}$. We have
$\Delta(L_{\alpha,k})$ $=$ $\Delta_{L_{0,i},L_{\alpha,k}}(L_{\alpha,k})$
$=$ $[a_{0,z}(L_{0,i},L_{\alpha,k})L_{0,z}+b_{0,z}(L_{0,i},L_{\alpha,k})H_{0,z},L_{\alpha,k}]+(1-\delta_{0,0})\lambda_{1}(L_{0,i},L_{\alpha,k})D_{f}(L_{\alpha,k})$
$+(1-\delta_{0,0})\lambda_{2}(L_{0,i},L_{\alpha,k})D_{g}(L_{\alpha,k})+\lambda_{3}(L_{0,i},L_{\alpha,k})D_{b}(L_{\alpha,k})$
$+(1-\delta_{i,0})\lambda_{4}(L_{0,i},L_{\alpha,k})D^{\rho}(L_{\alpha,k})$
$=$ $-\alpha a_{0,z}(L_{0,i},L_{\alpha,k})L_{\alpha,z+k}+k(1-\delta_{i,0})\lambda_{4}(L_{0,i},L_{\alpha,k})L_{\alpha,k-1}$,
$\Delta(L_{\alpha,k})$ $=$ $\Delta_{L_{1,j},L_{\alpha,k}}(L_{\alpha,k})$
$=$ $[a_{0,z}(L_{1,j},L_{\alpha,k})L_{1,z}+b_{0,z}(L_{1,j},L_{\alpha,k})H_{0,z},L_{\alpha,k}]+(1-\delta_{1,0})\lambda_{1}(L_{1,j},L_{\alpha,k})D_{f}(L_{\alpha,k})$
$+(1-\delta_{1,0})\lambda_{2}(L_{1,j},L_{\alpha,k})D_{g}(L_{\alpha,k})+\lambda_{3}(L_{1,j},L_{\alpha,k})D_{b}(L_{\alpha,k})$
$+(1-\delta_{j,0})\lambda_{4}(L_{1,j},L_{\alpha,k})D^{\rho}(L_{\alpha,k})$
$=$ $(1-\alpha)a_{1,z}(L_{1,j},L_{\alpha,k})L_{\alpha+1,z+k}+f_{\alpha}\lambda_{1}(L_{1,j},L_{\alpha,k})L_{\alpha,k}+g_{\alpha}\lambda_{2}(L_{1,j},L_{\alpha,k})L_{\alpha,k}$
$+k(1-\delta_{j,0})\lambda_{4}(L_{1,j},L_{\alpha,k})L_{\alpha,k-1}$.
@@
$a_{0,z}(L_{0,i},L_{\alpha,k})=a_{1,z}(L_{1,j},L_{\alpha,k})=\lambda_{1}(L_{1,j},L_{\alpha,k})=\lambda_{2}(L_{1,j},L_{\alpha,k})=0,$
@@
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$ such that $\Delta(L_{\alpha,i})=0$. Then for any $x=\sum_{\begin{subarray}{c}\gamma\in\Gamma\\ k\in\mathbb{Z}\end{subarray}}(m_{\gamma,k}L_{\gamma,k}+n_{\gamma,k}H_{\gamma,k})\in\mathcal{L}(\Gamma)$,
$\Delta(x)=0.$

### v2

@@ 2. Derivations on $\mathcal{L}(\Gamma)$
For all $\alpha\in\Gamma$ and all $i\in\mathbb{Z}$, assume that
$D(H_{\alpha,i})=\phi_{\alpha,i}L_{\alpha,i}+\psi_{\alpha,i}H_{\alpha,i},\,\phi_{\alpha,i},\psi_{\alpha,i}\in\mathbb{F}[t,t^{-1}].$
On the one hand,
$D([L_{-\alpha,-i},H_{\alpha,i}])$ == $D(-\alpha H_{0,0})$
== $\alpha(\phi_{0,0}L_{0,0}+\psi_{0,0}H_{0,0})$.
@@
$D([L_{-\alpha,-i},H_{\alpha,i}])$ == $D(L_{-\alpha,-i}),H_{\alpha,i}]+[L_{-\alpha,-i},D(H_{\alpha,i})]$
== $f_{-\alpha}L_{-\alpha,-i}+g_{-\alpha}H_{-\alpha,-i},H_{\alpha,i}]+[L_{-\alpha,-i},\phi_{\alpha,i}L_{\alpha,i}+\psi_{\alpha,i}H_{\alpha,i}]$
== $\alpha f_{-\alpha}H_{0,0}-2\alpha\phi_{\alpha,i}L_{0,0}-\alpha\psi_{\alpha,i}H_{0,0}$.
Then we have $\alpha(2\phi_{\alpha,i}-\phi_{0,0})L_{0,0}+\alpha(f_{-\alpha}+\psi_{\alpha,i}-\psi_{0,0})H_{0,0}=0$, thus
$\phi_{\alpha,i}=\frac{1}{2}\phi_{0,0},\,f_{-\alpha}+\psi_{\alpha,i}=\psi_{0,0}$
for all $0\neq\alpha\in\Gamma.$ Since $H_{0,0}\in C(\mathcal{L}(\Gamma))$, then $D(H_{0,0})\in C(\mathcal{L}(\Gamma))$. Thus $\phi_{0,0}=0$. Furthermore, $\phi_{\alpha,i}=0$ for all $\alpha\in\Gamma$ and all $i\in\mathbb{Z}$, we have
$D(H_{\alpha,i})$ == $\psi_{\alpha,i}H_{\alpha,i}$
== $(-f_{-\alpha}+\psi_{0,0})H_{\alpha,i}$
== $(f_{\alpha}+\psi_{0,0})H_{\alpha,i}$.
@@
$\mathrm{Der}\mathcal{L}(\Gamma)=\mathrm{ad}(\mathcal{L}(\Gamma))+(\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}])}\oplus\mathcal{D}_{\mathbf{g}(\Gamma)}\oplus\mathcal{D}_{\mathbb{F}[t,t^{-1}]}\oplus\mathcal{D}_{\mathbb{F}[t,t^{-1}]\frac{d}{dt}})$.
@@
(i) First we show that $(\mathrm{Der}\mathcal{L}(\Gamma))_{0}=\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}])}+\mathcal{D}_{\mathbf{g}(\Gamma)}+\mathcal{D}_{\mathbb{F}[t,t^{-1}]}+\mathcal{D}_{\mathbb{F}[t,t^{-1}]\frac{d}{dt}}$.
In fact, by Lemma 3 and Lemma 4, on one hand, take any $D\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$, for all $D_{\phi}\in\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}])},\,D_{g}\in\mathcal{D}_{\mathbf{g}(\Gamma)},\,D_{b}\in\mathcal{D}_{\mathbb{F}[t,t^{-1}]},\,D^{\rho}\in\mathcal{D}_{\mathbb{F}[t,t^{-1}]\frac{d}{dt}}$, such that
$(D-D^{\rho})(L_{\alpha,i})$ == $f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i}$,
$(D-D^{\rho})(H_{\alpha,i})$ == $(f_{\alpha}+b)H_{\alpha,i}$.
On the other hand, it follows that
$(D_{\phi}+D_{g}+D_{b})(L_{\alpha,i})$ == $f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i}$,
$(D_{\phi}+D_{g}+D_{b})(H_{\alpha,i})$ == $(f_{\alpha}+b)H_{\alpha,i}$.
Thus $D=D_{\phi}+D_{g}+D_{b}+D^{\rho}$.
(ii) Next we show that $\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}])}+\mathcal{D}_{\mathbf{g}(\Gamma)}+\mathcal{D}_{\mathbb{F}[t,t^{-1}]}+\mathcal{D}_{\mathbb{F}[t,t^{-1}]\frac{d}{dt}}$ is a direct sum.
@@
$D=D_{\phi}+D_{g}+D_{b}+D^{\rho},$
where $\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}]),\,g\in\mathbf{g}(\Gamma),\,b\in\mathbb{F}[t,t^{-1}],\,\rho\in\mathbb{F}[t,t^{-1}]\frac{d}{dt}$. If $D=0$, to prove that $D_{\phi}=D_{g}=D_{b}=D^{\rho}=0$. In fact, we have
$D(L_{\alpha,i})=0,$
$D(H_{\alpha,i})=0.$
@@
$(D_{\phi}+D_{g}+D_{b}+D^{\rho})(L_{\alpha,i})=\phi(\alpha)L_{\alpha,i}+g_{\alpha}H_{\alpha,i}+L_{\alpha}\rho(t^{i}),$
$(D_{\phi}+D_{g}+D_{b}+D^{\rho})(H_{\alpha,i})=\phi(\alpha)H_{\alpha,i}+bH_{\alpha,i}+H_{\alpha}\rho(t^{i}).$
Then $D_{\phi}=D_{g}=D_{b}=D^{\rho}=0$. ∎
@@ 3. $2$-local derivations on $\mathcal{L}(\Gamma)$
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$. For all $x,y\in\mathcal{L}(\Gamma)$, there exist derivations $\Delta_{x,y}$ on $\mathcal{L}(\Gamma)$ such that $\Delta(x)=\Delta_{x,y}(x)\,,\,\Delta(y)=\Delta_{x,y}(y)$ and it can be written as
$\Delta_{x,y}=\mathrm{ad}(\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ i\in\mathbb{Z}\end{subarray}}(a_{\alpha,i}(x,y)L_{\alpha,i}+b_{\alpha,i}(x,y)H_{\alpha,i}))+\lambda_{1}(x,y)D_{\phi}+\lambda_{2}(x,y)D_{g}+\lambda_{3}(x,y)D_{b}+\lambda_{4}(x,y)D^{\rho}.$
where $a_{\alpha,i},b_{\alpha,i},\,\lambda_{i}$ for $i=1,2,3,4$ are complex-valued functions on $\mathcal{L}(\Gamma)\times\mathcal{L}(\Gamma)$, $D_{\phi},D_{g},D_{b},D^{\rho}$ are given as previously.
@@
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$. For any fixed $x\in\mathcal{L}(\Gamma)$, if $\Delta(L_{\alpha,i})=0$, $\alpha\in\Gamma,i\in\mathbb{Z}$, then
$\Delta_{L_{\alpha,i},x}$ == $\mathrm{ad}(\sum_{\begin{subarray}{c}\alpha\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}(a_{\alpha,j}(L_{\alpha,i},x)L_{\alpha,j}+b_{0,j}(L_{\alpha,i},x)H_{0,j}))+(1-\delta_{\alpha,0})\lambda_{1}(L_{\alpha,i},x)D_{\phi}$
$(1-\delta_{\alpha,0})\lambda_{2}(L_{\alpha,i},x)D_{g}+\lambda_{3}(L_{\alpha,i},x)D_{b}+(1-\delta_{i,0})\lambda_{4}(L_{\alpha,i},x)D^{\rho}$.
@@
If $\Delta(L_{\alpha,i})=0$, then we have
$\Delta(L_{\alpha,i})$ == $\Delta_{L_{\alpha,i},x}(L_{\alpha,i})$
== $\sum_{\begin{subarray}{c}\beta\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}(a_{\beta,j}(L_{\alpha,i},x)L_{\beta,j}+b_{\beta,j}(L_{\alpha,i},x)H_{\beta,j}),L_{\alpha,i}]+\lambda_{1}(L_{\alpha,i},x)D_{\phi}(L_{\alpha,i})$
$\lambda_{2}(L_{\alpha,i},x)D_{g}(L_{\alpha,i})+\lambda_{3}(L_{\alpha,i},x)D_{b}(L_{\alpha,i})+\lambda_{4}(L_{\alpha,i},x)D^{\rho}(L_{\alpha,i})$
== $\sum_{\begin{subarray}{c}\beta\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}((\beta-\alpha)a_{\beta,j}(L_{\alpha,i},x)L_{\alpha+\beta,i+j}+\beta b_{\beta,j}(L_{\alpha,i},x)H_{\alpha+\beta,i+j})$
$\phi(\alpha)\lambda_{1}(L_{\alpha,i},x)L_{\alpha,i}+g_{\alpha}\lambda_{2}(L_{\alpha,i},x)L_{\alpha,i}+i\lambda_{4}(L_{\alpha,i},x)L_{\alpha,i-1}$
== 0.
Then $a_{\beta,j}(L_{\alpha,i},x)=0$ for $\alpha\neq\beta$, $b_{\beta,j}(L_{\alpha,i},x)=0$ for all $\beta\neq 0$, $\lambda_{1}(L_{\alpha,i},x)=0$ for $\alpha\neq 0$, $\lambda_{2}(L_{\alpha,i},x)=0$ for $\alpha\neq 0$, $\lambda_{4}(L_{\alpha,i},x)=0$ for $i\neq 0.$ Thus, the conclusion holds. ∎
@@
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$ such that $\Delta(L_{0,i})=\Delta(L_{1,j})=0$. Then $\Delta(L_{\alpha,k})=0,$ for all $\alpha\in\Gamma$ and all $k\in\mathbb{Z}$.
@@
Since $\Delta(L_{0,i})=\Delta(L_{1,j})=0$, then we can assume that
$\Delta_{L_{\gamma,z},x}$ == $\mathrm{ad}(a_{\gamma,l}(L_{\gamma,z},x)L_{\gamma,l}+b_{0,l}(L_{\gamma,z},x)H_{0,l})+(1-\delta_{\gamma,0})\lambda_{1}(L_{\gamma,z},x)D_{\phi}$
$(1-\delta_{\gamma,0})\lambda_{2}(L_{\gamma,z},x)D_{g}+\lambda_{3}(L_{\gamma,z},x)D_{b}+(1-\delta_{z,0})\lambda_{4}(L_{\gamma,z},x)D^{\rho}$,
where $L_{\gamma,z}=\{L_{0,i},L_{1,j}\},\,x\in\mathcal{L}(\Gamma)$. Let $x=L_{\alpha,k}$. Then we have
$\Delta(L_{\alpha,k})$ == $\Delta_{L_{0,i},L_{\alpha,k}}(L_{\alpha,k})$
== $a_{0,l}(L_{0,i},L_{\alpha,k})L_{0,l}+b_{0,l}(L_{0,i},L_{\alpha,k})H_{0,l},L_{\alpha,k}]+(1-\delta_{0,0})\lambda_{1}(L_{0,i},L_{\alpha,k})D_{\phi}(L_{\alpha,k})$
$(1-\delta_{0,0})\lambda_{2}(L_{0,i},L_{\alpha,k})D_{g}(L_{\alpha,k})+\lambda_{3}(L_{0,i},L_{\alpha,k})D_{b}(L_{\alpha,k})$
$(1-\delta_{i,0})\lambda_{4}(L_{0,i},L_{\alpha,k})D^{\rho}(L_{\alpha,k})$
== $\alpha a_{0,l}(L_{0,i},L_{\alpha,k})L_{\alpha,l+k}+k(1-\delta_{i,0})\lambda_{4}(L_{0,i},L_{\alpha,k})L_{\alpha,k-1}$,
$\Delta(L_{\alpha,k})$ == $\Delta_{L_{1,j},L_{\alpha,k}}(L_{\alpha,k})$
== $a_{1,l}(L_{1,j},L_{\alpha,k})L_{1,l}+b_{1,l}(L_{1,j},L_{\alpha,k})H_{0,l},L_{\alpha,k}]+(1-\delta_{1,0})\lambda_{1}(L_{1,j},L_{\alpha,k})D_{\phi}(L_{\alpha,k})$
$(1-\delta_{1,0})\lambda_{2}(L_{1,j},L_{\alpha,k})D_{g}(L_{\alpha,k})+\lambda_{3}(L_{1,j},L_{\alpha,k})D_{b}(L_{\alpha,k})$
$(1-\delta_{j,0})\lambda_{4}(L_{1,j},L_{\alpha,k})D^{\rho}(L_{\alpha,k})$
== $(1-\alpha)a_{1,l}(L_{1,j},L_{\alpha,k})L_{\alpha+1,l+k}+\phi(\alpha)\lambda_{1}(L_{1,j},L_{\alpha,k})L_{\alpha,k}+g_{\alpha}\lambda_{2}(L_{1,j},L_{\alpha,k})L_{\alpha,k}$
$k(1-\delta_{j,0})\lambda_{4}(L_{1,j},L_{\alpha,k})L_{\alpha,k-1}$.
@@
$a_{0,l}(L_{0,i},L_{\alpha,k})=a_{1,l}(L_{1,j},L_{\alpha,k})=\lambda_{1}(L_{1,j},L_{\alpha,k})=\lambda_{2}(L_{1,j},L_{\alpha,k})=0,$
@@
Let $\Delta$ be a $2$-local derivation on $\mathcal{L}(\Gamma)$ such that $\Delta(L_{\alpha,i})=0$, $\alpha\in\Gamma,i\in\mathbb{Z}$. Then
$\Delta(x)=0$
for any $x=\sum_{\begin{subarray}{c}\gamma\in\Gamma\\ k\in\mathbb{Z}\end{subarray}}(m_{\gamma,k}L_{\gamma,k}+n_{\gamma,k}H_{\gamma,k})\in\mathcal{L}(\Gamma)$.